OpenStudy (anonymous):

A wheel has a radius of 15 cm. Approximately how far does it travel in 4 revolutions?

OpenStudy (anonymous):

\[r=15\] so circumference is\[ 2\pi r\]=\[2\pi*15 = 30\] in 4 revolutions the wheel will travel \[4*30\pi = 120\]

OpenStudy (anonymous):

sorry \[120 \pi\]